Two point charges are located on the positive x-axis of a co-ordinate system. Charge is 2.0cm from the origin, another charge is 4.0cm from the origin. What is the total force exerted by these two charges on a charge located at the origin?

1
Expert's answer
2022-02-09T13:39:51-0500

Solution

Let two charges are Q is at d=0.02m and Q' is at d'0.04m.

And let Q" Is at origin.

So net force on charge Q" Due to charges Q' and Q" is given by

"F=\frac{KQQ"}{d}+\\frac{KQ'Q"}{d'}\\=KQ"(\\frac{Q}{0.02}+\\frac{Q'}{0.04}) \\\\=25KQ"(2Q+Q')"

So this is force.